OpenStudy (harleygirl18):

can anyone check my answer choice Muriel opened a savings account with an annual interest rate of 9% and an initial deposit of $4000. If her interest is compounded quarterly, how much is in her account after 6 years? Round your answer to the nearest cent. interest compounded quarterly: A = P (1 +r/4)4t Answer:98160

OpenStudy (danjs):

t is in years P is the starting amount r is the annual rate decimal 0.09

OpenStudy (danjs):

each quarter of the year, the current value will be increased by (9/4)%, so each year you are multiplying the principal by 1.0225 , 4 times general exponential y=A*b^x

OpenStudy (harleygirl18):

so it would equal 31644.33

OpenStudy (danjs):

A = P *(1 +r/4)^(4t) \[\huge A = 4000*(1+\frac{ 0.09 }{ 4 })^{4*6}\]

OpenStudy (danjs):

there are 24 compound times 4*6=24, and each time it adds a factor of (1+.09/4)=1.0225 multiplied on again

OpenStudy (danjs):

i think it comes to near 6823 and change

OpenStudy (harleygirl18):

thank you sooooo much I've been stuck on this question for hours

OpenStudy (danjs):

no problem, it is just an exponential function y=A*b^x just have to get the compounding period and things down